# A functional central limit theorem for weighted occupancy processes of the Karlin model

### v2

@@ 3.1. Finite-dimensional convergence for the Poissonized model
$\sum_{\ell=\ell_{n}}^{\infty}{\mathbb{E}}\left(a_{N_{\ell}(ns)}a_{N_{\ell}(nt)}\right)$ $=\sum_{\ell=\ell_{n}}^{\infty}\sum_{j=1}^{\infty}\sum_{j^{\prime}=0}^{\infty}a_{j}a_{j+j^{\prime}}\frac{(nsp_{\ell})^{j}(n(t-s)p_{\ell})^{j^{\prime}}}{j!j^{\prime}!}e^{-ntp_{\ell}}$
$=\sum_{\ell=\ell_{n}}^{\infty}\sum_{k=1}^{\infty}\sum_{j=1}^{k}a_{j}a_{k}\frac{(nsp_{\ell})^{j}(n(t-s)p_{\ell})^{k-j}}{j!(k-j)!}e^{-ntp_{\ell}}$
$=\sum_{\ell=\ell_{n}}^{\infty}\sum_{k=1}^{\infty}\frac{F_{\vec{a},s,t}^{(1)}(k)(ntp_{\ell})^{k}}{k!}e^{-ntp_{\ell}}=\sum_{\ell=\ell_{n}}^{\infty}{\mathbb{E}}F_{\vec{a},s,t}^{(1)}(N_{\ell}(nt)).$
@@
$\sum_{\ell=\ell_{n}}^{\infty}{\mathbb{E}}\left(a_{N_{\ell}(ns)}a_{N_{\ell}(nt)}\right)\sim\sigma_{n}^{2}\alpha t^{\alpha}\sum_{k=1}^{\infty}\frac{F_{\vec{a},s,t}^{(1)}(k)\gamma(k-\alpha,t\epsilon^{-1/\alpha})}{k!}.$
@@
$\sum_{\ell=\ell_{n}}^{\infty}{\mathbb{E}}a_{N_{\ell}(ns)}{\mathbb{E}}a_{N_{\ell}(nt)}$ $=\sum_{\ell=\ell_{n}}^{\infty}\sum_{j=1}^{\infty}\sum_{j^{\prime}=1}^{\infty}a_{j}a_{j^{\prime}}\frac{(nsp_{\ell})^{j}(ntp_{\ell})^{j^{\prime}}}{j!j^{\prime}!}e^{-n(s+t)p_{\ell}}$
$=\sum_{\ell=\ell_{n}}^{\infty}\sum_{k=1}^{\infty}F_{\vec{a},s,t}^{(2)}(k)\frac{(n(s+t)p_{\ell})^{k}}{k!}e^{-n(s+t)p_{\ell}}$
@@
Next, we apply the central limit theorem for an infinite triangular array of independent random variables satisfying the Lyapunov-type condition ([9, Theorem 8.4.1]). Note that
$\sum_{i=1}^{d}c_{i}\widetilde{W}_{n}^{(\epsilon)}(t_{i})=\sum_{\ell=\ell_{n}}^{\infty}\sum_{i=1}^{d}c_{i}\left(a_{N_{\ell}(nt_{i})}-{\mathbb{E}}a_{N_{\ell}(nt_{i})}\right)$
is a summation of independent random variables, and we have verified the convergence of the variance as desired in (3.6). It remains to verify
(3.7) $\lim_{n\to\infty}\frac{1}{\sigma_{n}^{2+\delta}}\sum_{\ell=\ell_{n}}^{\infty}{\mathbb{E}}\left|\sum_{i=1}^{d}c_{i}\left(a_{N_{\ell}(nt_{i})}-\mathbb{E}a_{N_{\ell}(nt_{i})}\right)\right|^{2+\delta}=0,$
@@
$\frac{1}{\sigma_{n}^{2+\delta}}\sum_{\ell=\ell_{n}}^{\infty}{\mathbb{E}}\left|\sum_{i=1}^{d}c_{i}\left(a_{N_{\ell}(nt_{i})}-\mathbb{E}a_{N_{\ell}(nt_{i})}\right)\right|^{2+\delta}$ $\leq\frac{C}{\sigma_{n}^{2+\delta}}\sum_{\ell=\ell_{n}}^{\infty}\sum_{i=1}^{d}{\mathbb{E}}\left|a_{N_{\ell}(nt_{i})}-\mathbb{E}a_{N_{\ell}(nt_{i})}\right|^{2+\delta}$
$\leq\frac{C}{\sigma_{n}^{2+\delta}}\sum_{\ell=\ell_{n}}^{\infty}\sum_{i=1}^{d}{\mathbb{E}}|a_{N_{\ell}(nt_{i})}|^{2+\delta}\leq\frac{C\sigma_{n}^{2}}{\sigma_{n}^{2+\delta}}\to 0,$
@@
(3.8) $\mathbb{E}|X-{\mathbb{E}}X|^{2+\delta}\leq 2^{1+\delta}(\mathbb{E}|X|^{2+\delta}+|\mathbb{E}X|^{2+\delta})\leq 2^{\delta+2}\mathbb{E}|X|^{2+\delta},$
and in third inequality we used Lemma 3.5 under the assumption $|b_{j}|=|a_{j}|^{2+\delta}\leq Cj^{\alpha^{\prime}}$ for some $\alpha^{\prime}<\alpha$: for this it suffices to take $\delta>0$ such that $\alpha^{\prime}=\beta(2+\delta)<\alpha$. Therefore, (3.7) is verified.∎
@@
$\mathop{\rm Var}\left(\widetilde{T}_{n}^{(\epsilon)}(t)\right)\leq\sum_{\ell=1}^{\ell_{n}-1}\mathbb{E}(a_{N_{\ell}(nt)}^{2})\leq C\epsilon^{1-{2\beta}/{\alpha}}n^{\alpha}t^{2\beta}.$
@@ 3.2. Tightness for the Poissonized model
By Rosenthal’s inequality, we have
@@
The claimed tightness now follows from Lemma 3.9 below, which is an adaption of [15, Lemma 3.6]. Therein, the lemma assumes a single process $G(t)$ and our lemma here is a generalization in the sense that if the process $G_{n}(t)$ takes the form $G_{n}(t)=G(nt)$, then our Lemma 3.9 becomes [15, Lemma 3.6]. However, here we shall need $G_{n}(t)=\widehat{W}_{n}(t)$, and hence we need to slightly modify the tightness argument therein. We also found and filled in a small gap in the argument therein when establishing the corresponding (3.18) and (3.19) here; see comments below. ∎
@@
For every $p\in(0,\alpha/\beta)$, there exists $C>0$ such that
$\sum_{\ell=1}^{\infty}{\mathbb{E}}\left(N_{\ell}(nt)^{\beta p}{{\bf 1}}_{\left\{N_{\ell}(nt)\leq n/\log n\right\}}\right)\leq C\sigma_{n}^{2}t^{\alpha}\quad\mbox{ for all }n\in\mathbb{N},t\in[0,1].$
The upper bound is to be compared with (3.2) and (3.3) where the upper bound for the above without the indicator function was obtained. Without the indicator function, the upper bound in (3.2) is of the same order while in (3.3) we have obtained an upper bound of $C\sigma_{n}^{2}t^{\beta p}$. The gain here is we replace $t^{\beta p}$ by $t^{\alpha}$ (recall $p\beta<\alpha$).
@@
$\sum_{\ell=1}^{\infty}{\mathbb{E}}\left(N_{\ell}(nt)^{\beta p}{{\bf 1}}_{\left\{N_{\ell}(nt)\leq n/\log n\right\}}\right)$ $=\int_{0}^{\infty}\sum_{j=1}^{j_{n}}\frac{j^{\beta p}e^{-{nt}/{x}}}{j!}\left(\frac{nt}{x}\right)^{j}v({\rm d}x)$
$=\int_{0}^{\infty}\sum_{j=1}^{{j_{n}}}\frac{j^{\beta p}e^{-z}(j-z)z^{j-1}}{j!}v\left(\frac{nt}{z}\right){\rm d}z,$
where in the second step we used integration by parts and the substitution $z=nt/x$. Modifying the argument for (3.4) in the proof of Lemma 3.5, we can show
$\int_{0}^{1}\sum_{j=1}^{j_{n}}\frac{j^{\beta p}e^{-z}(j-z)z^{j-1}}{j!}v\left(\frac{nt}{z}\right){\rm d}z\leq C\sigma_{n}^{2}t^{\alpha}\quad\mbox{ for all }n\in\mathbb{N},t\in[0,1].$
For the integral over $[1,\infty)$, we first mention an identity: for any sequence $(D_{j})_{j\geq 0}$ with $D_{0}=0$ and $j_{n}\in\mathbb{N}$,
@@
Then, it follows that for $\delta\in(0,\alpha-p\beta)$ and constants $C$ depending on $\delta$,
$\int_{1}^{\infty}\sum_{j=1}^{j_{n}}\frac{j^{\beta p}e^{-z}(j-z)z^{j-1}}{j!}v\left(\frac{nt}{z}\right){\rm d}z$ $\leq v(nt)\int_{1}^{\infty}\sum_{j=0}^{j_{n}}{\frac{(j+1)^{\beta p}-j^{\beta p}}{j!}z^{j}}e^{-z}\frac{v(nt/z)}{v(nt)}{\rm d}z$
$\leq Cv(nt)\int_{1}^{\infty}\sum_{j=0}^{j_{n}}\frac{(j+1)^{\beta p}-j^{\beta p}}{j!}z^{j-\alpha+\delta}e^{-z}{\rm d}z,$
where with $D_{j}=j^{\beta p}$ we applied (3.13) in the first step (and removed the last negative term) and then applied Potter’s bound in the second step (here we needed the key property that $D_{j+1}-D_{j}\geq 0$). It remains to show that the integral above is bounded by a constant that does not depending on $n$. Indeed,
$\int_{1}^{\infty}$ $\sum_{j=0}^{j_{n}}\frac{(j+1)^{\beta p}-j^{\beta p}}{j!}z^{j-\alpha+\delta}e^{-z}{\rm d}z$
$=\int_{1}^{\infty}\sum_{j=1}^{j_{n}}\frac{j^{\beta p}(j-z)e^{-z}z^{j-1-\alpha+\delta}}{j!}{\rm d}z+\frac{(j_{n}+1)^{\beta p}}{j_{n}!}\int_{1}^{\infty}z^{j_{n}-\alpha+\delta}e^{-z}{\rm d}z$
$\leq\sum_{j=1}^{\infty}\frac{j^{\beta p}}{j!}(\alpha-\delta)\Gamma(j-\alpha+\delta)+\frac{(j_{n}+1)^{\beta p}\Gamma(j_{n}+1-\alpha+\delta)}{\Gamma(j_{n}+1)},$
which is finite since $\beta p+\delta<\alpha$. (Note that in the last step above for the first term, we first exchanged the integral and summation, next bounded $\sum_{j=1}^{j_{n}}\int_{1}^{\infty}$ by $\sum_{j=1}^{\infty}\int_{0}^{\infty}$, and then applied $j\Gamma(j-\alpha+\delta)-\Gamma(j+1-\alpha+\delta)=(\alpha-\delta)\Gamma(j-\alpha+\delta)$; if we first extend the summation and integral and then exchange the order, then Fubini’s theorem does not apply.) ∎
@@
    For some $p>\alpha$ we have
@@
(In [15, Proof of Lemma 3.6], the analysis of case where $s=t_{i}$ was missing, and it can be dealt with as shown here.) The expression of (3.19) is zero. Indeed, we have
@@ 4. De-Poissonization
The goal is to prove the convergence of $\sigma_{n}^{-1}(W_{n}(t))_{t\in[0,1]}$ and we have just proved the convergence of its Poissonized version $\widetilde{W}_{n}$ in Theorem 3.1. In this section we complete the proof of Theorem 1.1 by showing that the difference of the two processes is negligible in an appropriate sense. First, the two processes can be coupled as follows. Let $(\tau_{k})_{k\in\mathbb{N}}$ denote the consecutive arrival times of the Poisson process $(N(t))_{t\geq 0}$ with $N(t):=\sum_{\ell=1}^{\infty}N_{\ell}(t)$ and $\tau_{0}=0$. Set
@@
for $\epsilon\in(0,\alpha/2)$. We shall show that $\sup_{t\in[0,n^{-(1-\epsilon)}]}\sum_{\ell=1}^{\infty}(|\widetilde{q}_{\ell}(nt)|+|q_{\ell}(nt)|+|\widetilde{q}_{\ell}(n\lambda_{n}(t))|)=o(\sigma_{n})$. To see this, we first notice
@@
This is because the supremum is achieved at $n\lambda_{n}(n^{-(1-\epsilon)})=\tau_{\lfloor n^{\epsilon}\rfloor}$, which is a Poisson random variable with parameter $\lfloor n^{\epsilon}\rfloor$. Then, to show $\sup_{t\in[0,n^{-(1-\epsilon)}]}\sum_{\ell=1}^{\infty}\widetilde{q}_{\ell}(n\lambda_{n}(t))/\sigma_{n}\to 0$ in probability it suffices to show
@@
To better illustrate the analysis, we first deal with $t=1$ (instead of considering the supremum over $t\in[n^{-(1-\epsilon)},1]$).
A crucial step is to apply a local central limit theorem controlling $|\widetilde{q}_{\ell}(n)-q_{\ell}(n)|=|{\mathbb{E}}a_{N(np_{\ell})}-{\mathbb{E}}a_{B_{np_{\ell}}}|$ and also $|\widetilde{q}_{\ell}(n+x)-\widetilde{q}_{\ell}(n)|=|{\mathbb{E}}a_{N((n+x)p_{\ell})}-{\mathbb{E}}a_{N(np_{\ell})}|$, where $N(\lambda)$ is a Poisson random variable with parameter $\lambda>0$ and $B_{n,p}$ is a binomial random variable with parameters $(n,p)$. This step is technical and is explained in Section 4.1.
@@
faster than any polynomial rate (uniformly for all $p>n^{\epsilon}$). Indeed, $|a_{j}|\leq Cj^{\beta}$ and $\mathbb{P}(N(np)\notin J_{np})\to 0$ decays at rate $e^{-C\log^{2}n}$ (by Bennett’s inequality), and one can show for $n$ large enough,
@@
$\leq C\sup_{|u-n|\leq d_{n}}\sum_{\ell=m_{n}}^{\infty}{\mathbb{E}}\left(|N_{\ell}(u)-N_{\ell}(n)|^{\beta}{{\bf 1}}_{\left\{N_{\ell}(u)-N_{\ell}(n)\neq 0\right\}}\right)$
$\leq C\sum_{\ell=m_{n}}^{\infty}{\mathbb{E}}\left(N(d_{n}p_{\ell})^{\beta}{{\bf 1}}_{\left\{N(d_{n}p_{\ell})\neq 0\right\}}\right).$
@@
$\sum_{\ell=m_{n}}^{\infty}\left|\widetilde{q}_{\ell}(n\lambda_{n})-\widetilde{q}_{\ell}(n)\right|{\bf 1}_{A_{n}}$ $\leq C\sum_{\ell=m_{n}}^{\infty}\mathbb{P}\left(N_{\ell}(d_{n})\neq 0\right)\leq Cn^{1/2}\log n\sum_{\ell=m_{n}}^{\infty}p_{\ell}$
@@
We first consider $\ell\leq m_{n}$. Recall that $K_{n,\ell}$ is a binomial random variable with parameters $n,p_{\ell}$ (denoted also by $B_{n,p_{\ell}}$). This time, by Lemma 4.4 we have
@@
Proposition 4.1 is much more involved than the corresponding step in earlier examples. In [15], this step is relatively straightforward (especially for the estimate of $I_{n,2}$) as with $a_{j}=1$ or $a_{j}={{\bf 1}}_{\left\{j\rm~{}is~{}odd\right\}}$, the function $\widetilde{q}_{\ell}$ and $q_{\ell}$ have explicit formulae of which one can derive a sharp control on the difference above quickly. See Lemmas 4.2 and 4.4 therein. We also notice in passing that in [10, Step 4 in the proof of Theorem 3], the statement $\mathbb{P}(Y_{n,k}^{*}(t)=Z_{n,k}^{*}\mid\Pi(n\tau)=\left\lfloor nt\right\rfloor)=1$ is incorrect (it is correct if $Y_{n,k}^{*}$ and $Z_{n,k}^{*}(\tau)$ both are replaced by the same statistic without centering). This small overlook is easy to fix therein as they considered the simple case $a_{j_{0}}=1$ for some prescribed $j_{0}$ and $a_{j}=0$ otherwise. It is also worth pointing out that they did not apply the chaining argument for the tightness of the Poissonized process, while they still needed certain moment estimates on the increments as in our (3.9).

### v3

@@ 3.1. Finite-dimensional convergence for the Poissonized model
$\sum_{\ell=\ell_{n}+1}^{\infty}{\mathbb{E}}\left(a_{N_{\ell}(ns)}a_{N_{\ell}(nt)}\right)$ $=\sum_{\ell=\ell_{n}+1}^{\infty}\sum_{j=1}^{\infty}\sum_{j^{\prime}=0}^{\infty}a_{j}a_{j+j^{\prime}}\frac{(nsp_{\ell})^{j}(n(t-s)p_{\ell})^{j^{\prime}}}{j!j^{\prime}!}e^{-ntp_{\ell}}$
$=\sum_{\ell=\ell_{n}+1}^{\infty}\sum_{k=1}^{\infty}\sum_{j=1}^{k}a_{j}a_{k}\frac{(nsp_{\ell})^{j}(n(t-s)p_{\ell})^{k-j}}{j!(k-j)!}e^{-ntp_{\ell}}$
$=\sum_{\ell=\ell_{n}+1}^{\infty}\sum_{k=1}^{\infty}\frac{F_{\vec{a},s,t}^{(1)}(k)(ntp_{\ell})^{k}}{k!}e^{-ntp_{\ell}}=\sum_{\ell=\ell_{n}+1}^{\infty}{\mathbb{E}}F_{\vec{a},s,t}^{(1)}(N_{\ell}(nt)).$
@@
$\sum_{\ell=\ell_{n}+1}^{\infty}{\mathbb{E}}\left(a_{N_{\ell}(ns)}a_{N_{\ell}(nt)}\right)\sim\sigma_{n}^{2}\alpha t^{\alpha}\sum_{k=1}^{\infty}\frac{F_{\vec{a},s,t}^{(1)}(k)\gamma(k-\alpha,t\epsilon^{-1/\alpha})}{k!}.$
@@
$\sum_{\ell=\ell_{n}+1}^{\infty}{\mathbb{E}}a_{N_{\ell}(ns)}{\mathbb{E}}a_{N_{\ell}(nt)}$ $=\sum_{\ell=\ell_{n}+1}^{\infty}\sum_{j=1}^{\infty}\sum_{j^{\prime}=1}^{\infty}a_{j}a_{j^{\prime}}\frac{(nsp_{\ell})^{j}(ntp_{\ell})^{j^{\prime}}}{j!j^{\prime}!}e^{-n(s+t)p_{\ell}}$
$=\sum_{\ell=\ell_{n}+1}^{\infty}\sum_{k=1}^{\infty}F_{\vec{a},s,t}^{(2)}(k)\frac{(n(s+t)p_{\ell})^{k}}{k!}e^{-n(s+t)p_{\ell}}$
@@
To complete the proof of the central limit theorem it remains to verify
(3.7) $\lim_{n\to\infty}\frac{1}{\sigma_{n}^{2+\delta}}\sum_{\ell=\ell_{n}+1}^{\infty}{\mathbb{E}}\left|\sum_{i=1}^{d}c_{i}\left(a_{N_{\ell}(nt_{i})}-\mathbb{E}a_{N_{\ell}(nt_{i})}\right)\right|^{2+\delta}=0,$
@@
$\frac{1}{\sigma_{n}^{2+\delta}}\sum_{\ell=\ell_{n}+1}^{\infty}{\mathbb{E}}\left|\sum_{i=1}^{d}c_{i}\left(a_{N_{\ell}(nt_{i})}-\mathbb{E}a_{N_{\ell}(nt_{i})}\right)\right|^{2+\delta}$ $\leq\frac{C}{\sigma_{n}^{2+\delta}}\sum_{\ell=\ell_{n}+1}^{\infty}\sum_{i=1}^{d}{\mathbb{E}}\left|a_{N_{\ell}(nt_{i})}-\mathbb{E}a_{N_{\ell}(nt_{i})}\right|^{2+\delta}$
$\leq\frac{C}{\sigma_{n}^{2+\delta}}\sum_{\ell=\ell_{n}+1}^{\infty}\sum_{i=1}^{d}{\mathbb{E}}|a_{N_{\ell}(nt_{i})}|^{2+\delta}\leq\frac{C\sigma_{n}^{2}}{\sigma_{n}^{2+\delta}}\to 0,$
@@
(3.8) $\mathbb{E}|X-{\mathbb{E}}X|^{2+\delta}\leq 2^{1+\delta}(\mathbb{E}|X|^{2+\delta}+|\mathbb{E}X|^{2+\delta})\leq 2^{2+\delta}\mathbb{E}|X|^{2+\delta},$
and in the third inequality we used Lemma 3.5 under the assumption $b_{j}=|a_{j}|^{2+\delta}\leq Cj^{\alpha^{\prime}}$ for some $\alpha^{\prime}<\alpha$: for this it suffices to take $\delta>0$ such that $\alpha^{\prime}=\beta(2+\delta)<\alpha$. Therefore, (3.7) is verified.∎
@@
$\mathop{\rm Var}\left(\widetilde{T}_{n}^{(\epsilon)}(t)\right)\leq\sum_{\ell=1}^{\ell_{n}}\mathbb{E}(a_{N_{\ell}(nt)}^{2})\leq C\epsilon^{1-{2\beta}/{\alpha}}n^{\alpha}t^{2\beta}.$
@@ 3.2. Tightness for the Poissonized model
By Rosenthal’s inequality (see [29]), we have
@@
The claimed tightness now follows from Lemma 3.9 below. ∎
@@
For every $r\in(0,\alpha/\beta)$, there exists $C>0$ such that
$\sum_{\ell=1}^{\infty}{\mathbb{E}}\left(N_{\ell}(nt)^{\beta r}{{\bf 1}}_{\left\{N_{\ell}(nt)\leq n/\log n\right\}}\right)\leq C\sigma_{n}^{2}t^{\alpha}\quad\mbox{ for all }n\in\mathbb{N},t\in[0,1].$
The upper bound is to be compared with (3.2) and (3.3) where the upper bound for the above without the indicator function was obtained. Without the indicator function, the upper bound in (3.2) is of the same order while in (3.3) we have obtained an upper bound of order $C\sigma_{n}^{2}t^{\beta r}$. The gain here is we replace $t^{\beta r}$ by $t^{\alpha}$ (recall $r\beta<\alpha$).
@@
$\sum_{\ell=1}^{\infty}{\mathbb{E}}\left(N_{\ell}(nt)^{\beta r}{{\bf 1}}_{\left\{N_{\ell}(nt)\leq n/\log n\right\}}\right)$ $=\int_{0}^{\infty}\sum_{j=1}^{j_{n}}\frac{j^{\beta r}e^{-{nt}/{x}}}{j!}\left(\frac{nt}{x}\right)^{j}v({\rm d}x)$
$=\int_{0}^{\infty}\sum_{j=1}^{{j_{n}}}\frac{j^{\beta r}e^{-z}(j-z)z^{j-1}}{j!}v\left(\frac{nt}{z}\right){\rm d}z,$
where in the second step we used integration by parts and the substitution $z=nt/x$. Modifying the argument for the asymptotics of $A_{n}$ in (3.4) in the proof of Lemma 3.5, we can show
$\int_{0}^{1}\sum_{j=1}^{j_{n}}\frac{j^{\beta r}e^{-z}(j-z)z^{j-1}}{j!}v\left(\frac{nt}{z}\right){\rm d}z\leq C\sigma_{n}^{2}t^{\alpha}\quad\mbox{ for all }n\in\mathbb{N},t\in[0,1].$
(Indeed, the left-hand side above takes the same form of $A_{n}$ with $b_{j}$ replaced by $j^{\beta r}$ and the domain of integral $[0,ntp_{\ell_{n}}]$ replaced by $[0,1]$. Recall that $\lim_{n\to\infty}ntp_{\ell_{n}}=t\epsilon^{-1/\alpha}$.) For the integral over $[1,\infty)$, we first mention an identity: for any sequence $(D_{j})_{j\geq 0}$ with $D_{0}=0$ and $j_{n}\in\mathbb{N}$,
@@
Then, it follows that for $\delta\in(0,\alpha-r\beta)$ and constant $C$ depending on $\delta$,
$\int_{1}^{\infty}\sum_{j=1}^{j_{n}}\frac{j^{\beta r}e^{-z}(j-z)z^{j-1}}{j!}v\left(\frac{nt}{z}\right){\rm d}z$ $\leq v(nt)\int_{1}^{\infty}\sum_{j=0}^{j_{n}}{\frac{(j+1)^{\beta r}-j^{\beta r}}{j!}z^{j}}e^{-z}\frac{v(nt/z)}{v(nt)}{\rm d}z$
$\leq Cv(nt)\int_{1}^{\infty}\sum_{j=0}^{j_{n}}\frac{(j+1)^{\beta r}-j^{\beta r}}{j!}z^{j-\alpha+\delta}e^{-z}{\rm d}z,$
where with $D_{j}=j^{\beta r}$ we applied (3.13) in the first step (and removed the last negative term) and then applied Potter’s bound in the second step (here we needed the key property that $D_{j+1}-D_{j}\geq 0$). It remains to show that the integral above is bounded by a constant that does not depend on $n$. Indeed,
$\int_{1}^{\infty}$ $\sum_{j=0}^{j_{n}}\frac{(j+1)^{\beta r}-j^{\beta r}}{j!}z^{j-\alpha+\delta}e^{-z}{\rm d}z$
$=\int_{1}^{\infty}\sum_{j=1}^{j_{n}}\frac{j^{\beta r}(j-z)e^{-z}z^{j-1-\alpha+\delta}}{j!}{\rm d}z+\frac{(j_{n}+1)^{\beta r}}{j_{n}!}\int_{1}^{\infty}z^{j_{n}-\alpha+\delta}e^{-z}{\rm d}z$
$\leq\sum_{j=1}^{\infty}\frac{j^{\beta r}}{j!}(\alpha-\delta)\Gamma(j-\alpha+\delta)+\frac{(j_{n}+1)^{\beta r}\Gamma(j_{n}+1-\alpha+\delta)}{\Gamma(j_{n}+1)},$
which is finite since $\beta r+\delta<\alpha$. (Note that in the last step above for the first term, we first exchanged the integral and summation, next bounded $\sum_{j=1}^{j_{n}}\int_{1}^{\infty}$ by $\sum_{j=1}^{\infty}\int_{0}^{\infty}$, and then applied $j\Gamma(j-\alpha+\delta)-\Gamma(j+1-\alpha+\delta)=(\alpha-\delta)\Gamma(j-\alpha+\delta)$; if we first extend the summation and integral and then exchange the order, then Fubini’s theorem does not apply.) ∎
@@
    For some $p>1/\alpha$ we have
@@
This lemma is an adaption of [17, Lemma 3.6]. Therein, the lemma assumes a single process $G(t)$ and our lemma here is a generalization in the sense that if the process $G_{n}(t)$ takes the form $G_{n}(t)=G(nt)$, then our Lemma 3.9 becomes [17, Lemma 3.6]. However, here we shall need $G_{n}(t)=\widehat{W}_{n}(t)$, and hence we need to slightly modify the tightness argument therein.
@@
(In [17, Proof of Lemma 3.6], the analysis of case where $s=t_{i}$ was missing, and it can be dealt with as shown here.) The expression of (3.19) is zero. Indeed, we have
@@ 4. De-Poissonization
The goal is to prove the convergence of $\sigma_{n}^{-1}(W_{n}(t))_{t\in[0,1]}$ and we have just proved the convergence of its Poissonized version $\sigma_{n}^{-1}\widetilde{W}_{n}$ in Theorem 3.1. In this section we complete the proof of Theorem 1.1 by showing that the difference of the two processes is negligible in an appropriate sense. First, the two processes can be coupled as follows. Let $(\tau_{k})_{k\in\mathbb{N}}$ denote the consecutive arrival times of the Poisson process $(N(t))_{t\geq 0}$ with $N(t):=\sum_{\ell=1}^{\infty}N_{\ell}(t)$ and $\tau_{0}=0$. Set
@@
for $\epsilon\in(0,\alpha/2)$. We shall show that $\sup_{t\in[0,n^{-(1-\epsilon)}]}\sum_{\ell=1}^{\infty}(|\widetilde{q}_{\ell}(nt)|+|q_{\ell}(nt)|+|\widetilde{q}_{\ell}(n\lambda_{n}(t))|)=o(\sigma_{n})$ in probability. To see this, we first notice
@@
This is because the supremum of interest above is achieved at $t=n^{-(1-\epsilon)}$ and equals $n\lambda_{n}(n^{-(1-\epsilon)})=\tau_{\lfloor n^{\epsilon}\rfloor}$, which is a Poisson random variable with parameter $\lfloor n^{\epsilon}\rfloor$. Then, to show $\sup_{t\in[0,n^{-(1-\epsilon)}]}\sum_{\ell=1}^{\infty}\widetilde{q}_{\ell}(n\lambda_{n}(t))/\sigma_{n}\to 0$ in probability it suffices to show
@@
To better illustrate the analysis, we first deal with $t=1$ (instead of considering the supremum over $t\in[n^{-(1-\epsilon)},1]$). A crucial step is to apply a local central limit theorem controlling $|\widetilde{q}_{\ell}(n)-q_{\ell}(n)|=|{\mathbb{E}}a_{N(np_{\ell})}-{\mathbb{E}}a_{B_{n,p_{\ell}}}|$ and also $|\widetilde{q}_{\ell}(n+x)-\widetilde{q}_{\ell}(n)|=|{\mathbb{E}}a_{N((n+x)p_{\ell})}-{\mathbb{E}}a_{N(np_{\ell})}|$, where $N(\lambda)$ is a Poisson random variable with parameter $\lambda>0$ and $B_{n,p}$ is a binomial random variable with parameters $(n,p)$. This step is technical and is explained in Section 4.1.
@@
faster than any polynomial rate (uniformly for all $p>n^{\epsilon}$). Indeed, $|a_{j}|\leq Cj^{\beta}$ and $\mathbb{P}(N(np)\notin J_{np})\to 0$ decays at rate $e^{-C\log^{2}n}$ (by Cramér–Chernoff bounds [11]), and one can show for $n$ large enough,
@@
$\leq C\sup_{|u-n|\leq d_{n}}\sum_{\ell=m_{n}}^{\infty}{\mathbb{E}}\left(|N_{\ell}(u)-N_{\ell}(n)|^{\beta}\right)$
$\leq C\sum_{\ell=m_{n}}^{\infty}{\mathbb{E}}\left(N(d_{n}p_{\ell})^{\beta}\right).$
@@
$\sum_{\ell=m_{n}}^{\infty}\left|\widetilde{q}_{\ell}(n\lambda_{n})-\widetilde{q}_{\ell}(n)\right|{\bf 1}_{A_{n}}$ $\leq C\sum_{\ell=m_{n}}^{\infty}{\mathbb{E}}\left(N(d_{n}p_{\ell})\right)=Cn^{1/2}\log n\sum_{\ell=m_{n}}^{\infty}p_{\ell}$
@@
We first consider $\ell\leq m_{n}$. Recall that $K_{n,\ell}$ is a binomial random variable with parameters $n,p_{\ell}$ (denoted also by $B_{n,p_{\ell}}$). This time, since $\ell\leq m_{n}$ implies $p_{\ell}\in[n^{\alpha^{\prime}/\alpha},1/2]$ we can apply Lemma 4.4. We have
@@
Proposition 4.1 is much more involved than the corresponding step in earlier examples. In [17], this step is relatively straightforward (especially for the estimate of $I_{n,2}$) as with $a_{j}=1$ or $a_{j}={{\bf 1}}_{\left\{j\rm~{}is~{}odd\right\}}$, the function $\widetilde{q}_{\ell}$ and $q_{\ell}$ have explicit formulae of which one can derive a sharp control on the difference above quickly. See Lemmas 4.2 and 4.4 therein. We also notice in passing that in [12, Step 4 in the proof of Theorem 3], the statement $\mathbb{P}(Y_{n,k}^{*}(t)=Z_{n,k}^{*}\mid\Pi(n\tau)=\left\lfloor nt\right\rfloor)=1$ is incorrect (it is correct if $Y_{n,k}^{*}$ and $Z_{n,k}^{*}(\tau)$ both are replaced by the same statistic without centering). This small overlook is easy to fix therein as they considered the simple case $a_{j_{0}}=1$ for some prescribed $j_{0}$ and $a_{j}=0$ otherwise. It is also worth pointing out that they did not apply the chaining argument for the tightness of the Poissonized process, while they still needed certain moment estimates on the increments as in our (3.9).